# Geometric Stability: The Missing Axis of Representations

### v1

@@ 9.3.1 Top 10 Most Stable Models
eva02_large_patch14 0.94 0.40 eva02_large_patch14 0.95 1.10
resnext101_32x8d 0.92 0.60 inception_v3 0.90 1.03
inception_v3 0.90 0.35 resnext101_32x8d 0.88 1.12
vit_large_clip 0.87 0.84 vit_large_clip 0.87 1.18
inception_v4 0.86 0.36 inception_v4 0.85 1.03
vit_base_clip 0.84 0.62 vit_base_clip 0.84 1.12
swin_large 0.82 0.99 tf_efficientnetv2_s 0.78 1.08
swinv2_tiny 0.82 0.44 regnety_032 0.77 1.02
tf_efficientnetv2_s 0.82 0.50 efficientnetv2_rw_m 0.77 1.15
vit_base32_clip 0.80 0.52 vit_base32_clip 0.77 1.06
@@
efficientnetv2_rw_m 0.96 resnext101_32x8d 0.87 eva02_large_patch14 0.98
swin_base 0.96 inception_v4 0.86 eva02_base_patch14 0.97
@@
convnext_base 0.96 eva02_large_patch14 0.81 vit_base_clip 0.96
inception_v3 0.96 efficientnetv2_rw_m 0.80 vit_base_dinov2 0.96
@@
eva02_large_patch14 0.92 0.89
eva02_base_patch14 0.90 1.10
regnety_032 0.89 0.81
inception_v4 0.88 2.10
regnety_064 0.87 0.85
vit_large_clip 0.86 0.81
efficientnet_b0 0.86 0.92
@@ 9.3.2 Bottom 10 Least Stable Models
vit_base_patch16.augreg_in21k 0.22 1.17
deit3_small_patch16 0.23 1.09
@@
deit3_base_patch16 0.25 1.17
vit_tiny_patch16.augreg 0.27 1.05
vit_small_patch14_dinov2 0.30 1.19
vit_base_patch16.augreg_ft 0.31 1.21
@@ 9.4 Composite Scores
swin_large 1.97 swin_large 1.87 beit_base 2.40
vit_large.augreg 1.68 vit_large.augreg 1.66 vit_base.augreg 2.31
resnext101_32x8d 1.60 resnext101_32x8d 1.37 swin_large 1.76
beit_base 1.59 beit_base 1.28 resnext101_32x8d 0.96
vit_base.augreg 1.19 vit_base.augreg 1.01 vit_base.augreg_ft 0.90
@@
resnext101_32x8d 2.86 regnety_032 1.45 resnext101_32x8d 1.97
wide_resnet101_2 1.01 vit_large.augreg 1.44 wide_resnet101_2 1.17
efficientnetv2_rw_m 0.90 beit_base 1.30 inception_v3 1.06
beit_base 0.80 vit_base.augreg_ft 0.83 efficientnetv2_rw_m 0.89
inception_v3 0.74 regnety_064 0.83 inception_v4 0.86
@@ 9.5 Outlier Detection
vit_giant_patch14_dinov2 LogME HIGH 5/6
vit_large_patch14_dinov2 LogME HIGH 4/6
eva02_base_patch14 Shesha-Var HIGH 5/6
beit_base_patch16 LEEP HIGH 6/6
vit_base_patch16.augreg_in21k LEEP HIGH 5/6
deit3_small_patch16 Shesha-FS LOW 5/6
vit_large_patch14_dinov2 Shesha-FS LOW 3/6
@@ 9.6 Family Ranking Consistency
Inception 1 1 1 1 2 1
CLIP 2 2 3 3 3 2
EVA 3 3 5 4 1 3
Swin 4 4 4 8 4 7
EfficientNetV2 5 5 2 5 5 5
ConvNeXt 6 6 2 10 6 12
RegNetY 7 7 6 6 4 4
DeiT 9 9 7 9 8 8
ResNet 8 10 10 11 9 9
ViT 10 11 11 11 7 10
DeiT3 12 13 8 13 13 11
DINOv2 13 13 13 9 1 11
@@ 10.6 Model Lists
Model Family Models
MiniLM sentence-transformers/paraphrase-MiniLM-L3-v2,
sentence-transformers/all-MiniLM-L6-v2,
sentence-transformers/paraphrase-MiniLM-L6-v2,
sentence-transformers/all-MiniLM-L12-v2,
sentence-transformers/multi-qa-MiniLM-L6-cos-v1
MPNet sentence-transformers/all-mpnet-base-v2,
sentence-transformers/paraphrase-mpnet-base-v2,
sentence-transformers/multi-qa-mpnet-base-cos-v1
DistilBERT sentence-transformers/distilbert-base-nli-mean-tokens
DistilRoBERTa sentence-transformers/all-distilroberta-v1,
sentence-transformers/paraphrase-distilroberta-base-v1
BERT sentence-transformers/bert-base-nli-mean-tokens
RoBERTa sentence-transformers/stsb-roberta-base,
sentence-transformers/nli-roberta-base-v2
ALBERT sentence-transformers/paraphrase-albert-small-v2
GTE thenlper/gte-small, thenlper/gte-base, thenlper/gte-large
E5 intfloat/e5-small-v2, intfloat/e5-base-v2, intfloat/e5-large-v2
BGE BAAI/bge-small-en-v1.5, BAAI/bge-base-en-v1.5, BAAI/bge-large-en-v1.5
SimCSE princeton-nlp/sup-simcse-bert-base-uncased,
princeton-nlp/unsup-simcse-bert-base-uncased
@@ 11.1 Experimental Design
  • Source domain: IMDB movie reviews (maas-etal-2011-learning) (2,000 samples for metric computation)
@@
  • Experiment 2 target: Yelp review polarity (zhangCharacterlevelConvolutionalNetworks2015) (full transfer)

### v2

@@ 9.3.1 Top 10 Most Stable Models
eva02_large 0.94 0.40 eva02_large 0.95 1.10
resnext101_32x8d 0.92 0.51 inception_v3 0.90 1.01
inception_v3 0.90 0.29 resnext101_32x8d 0.88 1.12
vit_large_clip 0.87 0.84 vit_large_clip 0.87 1.17
inception_v4 0.86 0.40 inception_v4 0.85 1.04
vit_base_clip 0.84 0.63 vit_base_clip 0.84 1.10
swin_large 0.82 0.99 tf_efficientnetv2_s 0.78 1.07
swinv2_tiny 0.82 0.46 regnety_032 0.77 1.03
tf_efficientnetv2_s 0.82 0.46 efficientnetv2_rw_m 0.77 1.15
vit_base32_clip 0.80 0.58 vit_base32_clip 0.77 1.09
@@
efficientnetv2_rw_m 0.96 resnext101_32x8d 0.87 eva02_large 0.98
swin_base 0.96 inception_v4 0.86 eva02_base 0.97
@@
convnext_base 0.96 eva02_large 0.81 vit_base_clip 0.96
inception_v3 0.96 efficientnetv2_rw_m 0.80 vit_base_dinov2 0.95
@@
eva02_large 0.92 0.99
eva02_base 0.90 1.01
regnety_032 0.89 0.89
inception_v4 0.88 1.40
regnety_064 0.87 0.90
vit_large_clip 0.86 0.89
efficientnet_b0 0.86 0.97
@@ 9.3.2 Bottom 10 Least Stable Models
vit_base.augreg_in21k 0.22 1.17
deit3_small 0.23 1.11
@@
deit3_base 0.25 1.17
vit_tiny.augreg_ft 0.27 1.00
vit_small_patch14_dinov2 0.30 1.13
vit_base.augreg_ft 0.31 1.21
@@ 9.4 Composite Scores
beit_base 1.21 beit_base 1.01 vit_base.augreg 1.13
swin_large 1.00 swin_large 0.87 beit_base 1.02
vit_large_clip 0.91 vit_large_clip 0.86 inception_v3 0.27
vit_giant_dinov2 0.72 convnextv2_base 0.55 efficientnetv2_rw_m 0.23
efficientnetv2_rw_m 0.54 efficientnetv2_rw_m 0.54 poolformer_m36 0.22
@@
inception_v3 1.33 beit_base 1.21 wide_resnet101_2 0.98
wide_resnet101_2 1.22 vit_giant_dinov2 1.06 densenet201 0.94
efficientnetv2_rw_m 0.96 vit_large_dinov2 0.71 resnet152 0.93
wide_resnet50_2 0.80 swin_large 0.62 resnet50.a2 0.93
resnext50_32x4d 0.77 vit_large_clip 0.60 inception_v3 0.91
@@ 9.5 Outlier Detection
beit_base_patch16_224.in22k_ft_in22k LEEP HIGH 5/6
eva02_base_patch14_224.mim_in22k Shesha_Var HIGH 5/6
vit_base_patch16_224.augreg_in21k LEEP HIGH 4/6
vit_giant_patch14_dinov2.lvd142m LogME HIGH 4/6
deit3_small_patch16_224.fb_in1k Shesha_FS LOW 4/6
vit_large_patch14_dinov2.lvd142m LogME HIGH 3/6
@@ 9.6 Family Ranking Consistency
Inception 1 1 1 1 4 2
CLIP 2 2 18 2 2 3
EVA02 3 3 22 4 1 1
EfficientNetV2 7 6 4 3 5 10
MobileNetV3 8 9 12 8 6 8
ResNetRS 13 13 9 6 8 5
EfficientNet 9 10 13 9 7 7
RegNetY 10 7 20 5 13 6
ResNeXt 5 4 17 10 11 15
Swin 4 5 10 15 12 16
ResNetV2 14 11 19 11 9 11
SwinV2 6 8 2 19 14 26
DINO 16 19 14 20 19 4
ConvNeXt 12 14 5 21 15 27
WideResNet 26 20 16 7 17 9
PVTv2 17 15 7 16 22 22
RegNetX 19 12 24 12 21 14
CoAtNet 15 18 3 26 16 28
BEiT 11 26 28 22 10 12
MaxViT 23 24 6 17 18 23
ConvNeXtV2 18 17 8 18 25 29
DenseNet 25 21 25 13 20 13
PoolFormer 21 23 11 28 26 17
MAE 20 16 27 14 28 21
DeiT 24 25 15 23 27 20
ResNet 22 22 23 25 24 18
DINOv2 29 28 29 24 3 25
ViT 27 27 26 27 23 19
DeiT3 28 29 21 29 29 24
@@ 10.6 Model Lists
Model Family
sentence-transformers/paraphrase-MiniLM-L3-v2 MiniLM
sentence-transformers/all-MiniLM-L6-v2 MiniLM
sentence-transformers/paraphrase-MiniLM-L6-v2 MiniLM
sentence-transformers/all-MiniLM-L12-v2 MiniLM
sentence-transformers/multi-qa-MiniLM-L6-cos-v1 MiniLM
sentence-transformers/all-mpnet-base-v2 MPNet
sentence-transformers/paraphrase-mpnet-base-v2 MPNet
sentence-transformers/multi-qa-mpnet-base-cos-v1 MPNet
sentence-transformers/distilbert-base-nli-mean-tokens DistilBERT
sentence-transformers/all-distilroberta-v1 DistilRoBERTa
sentence-transformers/paraphrase-distilroberta-base-v1 DistilRoBERTa
sentence-transformers/bert-base-nli-mean-tokens BERT
sentence-transformers/stsb-roberta-base RoBERTa
sentence-transformers/nli-roberta-base-v2 RoBERTa
sentence-transformers/paraphrase-albert-small-v2 ALBERT
thenlper/gte-small GTE
thenlper/gte-base GTE
thenlper/gte-large GTE
intfloat/e5-small-v2 E5
intfloat/e5-base-v2 E5
intfloat/e5-large-v2 E5
BAAI/bge-small-en-v1.5 BGE
BAAI/bge-base-en-v1.5 BGE
BAAI/bge-large-en-v1.5 BGE
princeton-nlp/sup-simcse-bert-base-uncased SimCSE
princeton-nlp/unsup-simcse-bert-base-uncased SimCSE
@@ 11.1 Experimental Design
  • Source domain: IMDB movie reviews (Maas et al.,, 2011) (2,000 samples for metric computation)
@@
  • Experiment 2 target: Yelp review polarity (Zhang et al.,, 2015) (full transfer)